OpenStudy (anonymous):

equation is f(t)=30(3)t predit when all Americans(both past and present)) will have heard the rumor. assume the u.s population is 305 million

OpenStudy (anonymous):

im assuming you meant 30(3^t)??

OpenStudy (anonymous):

in that case, set the function equal to 305 million divide both sides by 30 then, if you take ln of both sides, remember you can move the t from the exponent in front of the natural log. do this step divide both sides by ln(3)

OpenStudy (anonymous):

30(3^t)=305mil 3^t=305mil/30 ln(3^t)=ln(305mil/30) t*ln(3)=ln(305mil/30) t=ln(305mil/30)/ln(3)

OpenStudy (anonymous):

answer = 14.68

OpenStudy (anonymous):

i dont get that answer

OpenStudy (anonymous):

make sure you use parenthesis the right way 305 million = 305,000,000 t=ln(305000000/30)/ln(3)